OpenStudy (anonymous):

What is the slope of the line that contains the points (-1, 9) and (5, 21)? Enter your answer as an integer.

OpenStudy (tiffany_rhodes):

(21-9)/(5- -1) = 12/6 = 2. I just used the rise over run formula

OpenStudy (tiffany_rhodes):

(y2 - y1) /( x2 - x1) = m (the slope)